maths89898:

If s = 1 over 3 unit and A = 18s2, what is the value of A, in square units? ____ square unit(s). (Input whole number only.)

PDOfficerLaw:

@pdofficerlaw wrote:
"If s = 1 over 3 unit " \[1/3\] "A = 18s2" \[a=18(1/3)^2\] Multiply the squared first to 1/3. \[=18(1/9)\] Multiply 1/9 to 18 \[=2\] 2 squared units[/s]

maths89898:

tysm!

PDOfficerLaw:

No problem! :)

dukedadon:

the level 1s out here doing better than us LMAO